# Physics-aware Machine Learning Revolutionizes Scientific Paradigm for Machine Learning and Process-based Hydrology

### v1

@@ 3 HydroPML: Process-based Hydrology in Physics-aware Machine Learning
As shown in Fig. 5, the proposed HydroPML includes hydrodynamic process understanding and rainfall-runoff hydrological process understanding.
@@
Figure 6: The operational flood forecasting system based on HydroPML
@@ 3.1 Rainfall-runoff Process Understanding
Rainfall-runoff models are extensively utilized in hydrology to investigate hydrological processes and play a crucial role in water resources management, encompassing flood prediction [203] and drought analysis [204]. These models have been developed for diverse applications, ranging from small catchments to global scales.
3.1.1 Physical Methods in Hydrological Process
The hydrological model can be classified into three broad groups: the empirical model, conceptual models, and physical-based models [2]. Empirical models only take information from the existing data without considering the features and processes of hydrological systems and hence these models are also called data-driven models. Physical and conceptual models need a thorough understanding of the physics involved in the movement of surface water in the hydrological cycle [205].
Conceptual models interpret runoff processes by connecting simplified components in the overall hydrological process with the help of a number of interconnected storages used to represent different components. The conceptual models are usually lumped in nature and use the same value of parameters for the whole watershed. These models need a range of parameters and meteorological input data. A large number of meteorological and hydrological records is required for parameter calibration. A host of conceptual models have been proposed in the past including HBV model [206], non-recorded catchment areas (NRECA) [207], Boughton model [208], ARNO model [209], Xinanjiang model [210], etc.
Physical-based models are based on an understanding of the physics related to hydrological processes. The general physics laws and principles used include water balance equations, conservation of mass and energy, momentum, and kinematics. Finite difference or finite computation schemes are generally used to solve these equations. For being built on physical laws and equations, spatially-explicit, process-based models are, in theory, capable to respond correctly to dynamic changes in catchment conditions and climate forcing; they can thus be used for the assessment of non-linearities between states, drivers, and catchment responses [211, 212]. Therefore, they are practical tools to predict, project, and attribute impacts of global change (e.g. climate change, land use change, water management) on the water cycle and even hydrological extremes. The greatest strength of a physical model is the connection between model parameters and physical catchment characteristics, which makes it more realistic. As physical models are watershed semi- or fully distributed in nature, they require a large amount of data such as soil-hydraulic properties, vegetation, land cover characteristics, initial water depth, topography, or dimensions of the river network to explore the changes in the hydrological cycle due to human interference and climate change for water resources management [213]. Some examples of physical models include soil and water assessment tool (SWAT) [214], system hydrologique european (SHE) [215] or MIKE SHE [216], visualizing ecosystem land management assessments (VELMA) [217], kinematic runoff and erosion model (KINEROS) [218], Penn State integrated hydrologic modeling system (PIHM) [219], variable infiltration capacity (VIC) [220], and others.
@@ 3.1.2 Rainfall-runoff Forecast Meets Physics-aware Machine Learning
(1.1) Short-term Forecasts
LSTM [221, 222], physics-guided LSTM [43], GRUs [223], graph convolutional GRUs [224], graph convolutional RNN [225], stacked LSTM and bi-directional LSTM [226], ensemble model [227, 228], ConvLSTM [229], data-driven forecasting using only in-stream measurements [230]
GW-PINN [231]
@@
Hybrid method combining hydrologic processes with LSTM [232], SWAT-LSTM [233], conceptual hydrological model by integrating NNs [234], urban rainfall-runoff modeling [235], Google’s operational flood forecasting system [236]
(1.2) Long-term Forecasts
LSTM [10], Gaussian process regression [237], deep belief network [238], extreme learning machine [239], an encoder-decoder CNN algorithm [240], Bayesian ensemble learning by combining different ML methods [241]
LSTM by considering conservation of mass into loss function [15]
Entity-aware-LSTM [242], MC-LSTM [168], differentiable modeling [13], P-RNN layer [9]
A hybrid monthly streamflow forecasting approach [194], ML is embedded into a lumped conceptual rainfall-runoff model [195], three hybridization approaches [197], differentiable parameter learning [243]
(1.3) Spatial Variability
@@
Fully-distributed GNN [244], graph convolutional GRUs based model [224], graph convolutional RNN [225], ConvLSTM [229], graph-based reinforcement learning [245], data synergy [246], initializing a recurrent GNN by physical model [247]
HydroNets [248], a differentiable, learnable physics-based routing model [249]
Hybrid approach combining a spatially-distributed snow model with convLSTM [250]
(1.4) Bias Reduction and Reliability
Bias-corrected remote sensing products [251], directly measured datasets [230], ensemble methods [252, 253]
MC-LSTM [168, 254], differentiable modeling [13]
@@
LSTM using an additional input feature from the process-based model [255], using physical models coupled with LSTM [256], WRF-Hydro-LSTM [257], LSTM daily streamflow prediction models [258]
(1.5) Missing Data and Ungauged Basins
@@
Attention-based deep learning (such as 3D-CNN-Transformer) [259], input-selection ensemble method [260]
AI-based hybrid models [196], serial hybrid LSTMs [261, 262], transfer learning in SWAT-LSTM [233].
(1.6) Parameterization
@@
A machine learning approach [263], deep neural network [264], machine learning-based maps [265]
A differentiable neural network [202], differentiable parameter learning [243]
(1.7) Uncertainty Estimation in Rainfall-runoff Forecast
@@
Data-driven approach [266], Bayesian processor [267], hybrid ensemble and variational data assimilation framework [268], hierarchical model tree [269], probabilistic deep ensembles [270], LSTM-based approaches [271, 272], uncertainty estimation benchmarking procedure and baselines [273], analysis framework for sources of uncertainties [274]
Using physical models coupled with LSTM [256], XAJ-MCQRNN [275], hybrid error correction model [276]
@@
Short-term Forecasts. See Fig. 5 (label 1.1). Short-term forecasts focus on outlook horizons of hours to weeks. Innovative forecasting tools will reform river discharge forecasting, flood early warning mechanisms, and agricultural and hydropower management schemes. PaML offers robust tools to address the formidable challenges associated with the uncertainties, data dependencies, and intricacies inherent in short-term hydrological forecasting. First, the high uncertainty and complexity associated with hydrological processes and weather-climate factors affecting river basins have increasingly led researchers to use PDgML approaches. Specifically, the PDgML models based on LSTM [221, 222] are proposed to predict streamflow. They have been found to have greater accuracy in predicting hourly and daily streamflow than a classical neural network with back-propagation. A model [223] that uses multiple GRUs is proposed to predict up to 120 hours of streamflow. Granata et al. [227] show how the ensemble model based on random forest and multilayer perceptron outperforms the bi-directional LSTM network in predicting peaks of flow rates, as well as having significantly shorter computation times. Wegayehu and Muluneh [226] also compare different ML algorithms for one-step daily streamflow forecasting and show how both MLP and GRU algorithms outperform stacked LSTM and bi-directional LSTM on a nearly equal basis. An ensemble model [228] based on the combination of the nonlinear autoregressive network, multilayer perceptron, and random forest is proposed for short-term streamflow forecasting, considering precipitation as the only exogenous input and a forecast horizon of up to 7 days. In addition, to model the spatial diversity of rivers, the graph convolutional GRUs-based model [224] is developed to predict the next 36 hours of streamflow for a sensor location using the upstream river network. Graph convolutional RNN [225] is developed for robust water demand forecasting. More detailed information about PDgML studies on short-term streamflow prediction can be found in [277]. Furthermore, to overcome data dependence, attempts have been made based on the PiML community, such as PINNs, for solving groundwater flow equations [231]. Because it is difficult to express the whole process of rainfall-runoff with specific equations, the current progress based on PiML technology is very slow. In addition, the hybrid method combining hydrologic processes with LSTM [232] is successful in assimilating recent observations of streamflow to improve near-term daily forecasts of streamflow. A coupled soil and water assessment tool (SWAT)-LSTM approach [233] is developed to provide a potential shortcut to achieving daily streamflow simulations in both ungauged and poorly gauged watersheds. The conceptual hydrological model by integrating NNs [234] reveals that the hybrid model has better prediction performance than either the original GR4J or the single NN–based runoff prediction model.
@@
In general, existing physical and conceptual models are well-established in operational forecasting within the time scale of hours to weeks. The development of PaML-based short-term forecast methods is not very mature [30]. However, PaML offers promising alternatives to physics-based hydrological and inundation models for flood forecasting due to its automation, efficiency, and flexibility. For example, the PDgML (ConvLSTM [229], data-driven forecasting using only in-stream measurements [230]) for extracting spatio-temporal features of hydrological information outperforms the recent models in terms of flood arrival time and peak discharge. PaHL-based methods (such as Google’s operational flood forecasting system [236] and urban rainfall-runoff modeling [235]) show potential for flood nowcasting.
@@
Long-term Forecasts. See Fig. 5 (label 1.2). Long-term forecasts focus on outlook horizons of sub-seasonal, year, to decade. The vast majority of ML predictions in the literature are PDgML methods, which are driven by observational or historical data. PDgML allows us to capture complex non-linear patterns from data that would otherwise require extrapolation, thus providing an alternative approach to long-term rainfall-runoff forecasts, for example, LSTM [10], Gaussian process regression [237], deep belief network [238], extreme learning machine [239], an encoder-decoder CNN algorithm [240], and Bayesian ensemble learning by combining different ML methods [241]. These methods have good accuracy in the study of the rainfall-runoff relationship and climate factor-rainfall relationship. In order to alleviate the dependence on data, a physics-guided LSTM [15] that considers the conservation of mass into loss function is proposed. Digging deeper into physics (such as mass conservation and similarities) in the hydrological process, MC-LSTM [168] is proposed by modifying the LSTM to guarantee the conservation of the mass input. Entity-Aware-LSTM [242] is presented for learning catchment similarities as a feature layer in a deep learning model. In addition, based on the PeML technology, some NN modules with physical properties (such as differentiable modeling [13] and the P-RNN layer in the general physics-AI approach [9]) are proposed for rainfall-runoff forecasts. Furthermore, hybrid learning models (PaHL) can be suitable alternatives to hydrological models. Hybrid learning based on conceptual hydrological models tries to combine the strengths of hydrological models and data-driven methods. For example, a hybrid monthly streamflow forecasting approach [194] and three hybridization approaches [197] show that the hybrid model is not only significantly more precise than conceptual hydrological models but also outperforms a purely ML model.
@@
However, in most hybrid modeling frameworks, variables computed by the conceptual model are used as inputs to a data-driven model, which necessarily increases computation time. In addition, those interactions often go unaccounted for because the two models are calibrated separately. Thus, embedded ML in a lumped conceptual rainfall-runoff model [195] is proposed by coupling the two models and performing their calibration jointly, and differentiable parameter learning [243] is presented to efficiently learn a hybrid model in an end-to-end manner.
@@
Spatial Variability. See Fig. 5 (label 1.3). Rainfall-runoff modeling is difficult due to the non-linear structure of the process, which involves much spatio-temporal variability. Spatial variability is critical for predictions in ungauged basins and for developing universal regional models. PaML (especially GNNs) provides the possibility of distributed modeling. Specifically, for PDgML, GNN [244] is used for fully-distributed rainfall-runoff modeling. The result shows that the GNN-based fully-distributed model [244] successfully represents the spatial information in predictions. A graph convolutional GRUs based model [224], ConvLSTM [229], and graph convolutional RNN [225] can also be developed to model spatial diversity of rivers. A recurrent GNN [247] is proposed to capture the interactions among multiple segments in the river network. In this work, a physics-based model is used to initialize the machine learning model and learn the physics of streamflow. To reduce the need for large training datasets, a real-time active learning method that uses spatial and temporal contextual information to select representative query samples is employed in a graph-based reinforcement learning framework [245]. The effectiveness of graph-based reinforcement learning is demonstrated by predicting streamflow and water temperature in the Delaware River basin. In addition, region-specific differences can be learned for the LSTM model by inputting sufficient heterogeneous data (data synergy) [246].
@@
Furthermore, spatial physical characteristics of rivers (such as river networks and hydrologic routing modules) can be encoded into NNs based on the PeML technology. For example, a hydrologic model called HydroNets [248] is built by leveraging river network structure. Specifically, each node in the graph represents a basin, and an edge direction corresponds to water flow from a sub-basin to its containing basin. A differentiable, learnable physics-based routing model [249] is proposed to constrain spatially-distributed parameterization. This is the first time an interpretable, physics-based model is learned on the river network to infer spatially-distributed parameters. Finally, for the PaHL, the hybrid models combining conceptual models and data-driven models are often lumped, i.e., spatially averaged, at the catchment scale. To address this problem, a hybrid approach [250] is developed to predict streamflow. The hybrid approach combines a spatially-distributed snow model with convLSTM, which simulates the effect of surface and subsurface properties on the streamflow. In the future, building a robust spatio-temporal representation for different watersheds based on the PaML community will be an ongoing direction for research.
@@
Bias Reduction and Reliability. See Fig. 5 (label 1.4). One important challenge of PaML-based rainfall-runoff forecasts is the ability of such models to produce reliable or physically-plausible results in different domains, especially extreme conditions such as floods and droughts. Specifically, reliable spatio-temporal remote sensing data and directly measured datasets can be used to reduce the bias of PDgML-based rainfall-runoff forecasting. For example, bias-corrected remote-sensing precipitation products are adopted as precipitation input of a spatio-temporal DL rainfall-runoff forecasting model [251], which can improve the model’s reliability for extreme flood forecasts. Directly measured datasets are utilized for streamflow estimation to develop a data-driven forecasting algorithm [230]. However, these bias-corrected data-based approaches suffer from deficiencies over small catchments due to uncertainty in remote sensing data and errors in hydrological models. Thus, some ensemble methods (such as resampling ensemble methods [252], which combine hydrometeorological modeling and LSTM [253]) are explored to improve the reliability and robustness of PDgML models. Furthermore, one emerging route for bias reduction and reliability is to employ PeML designs that explicitly observe the physical laws in the rainfall-runoff process, such as MC-LSTM [168] and differentiable modeling [13]. Another new development is hybrid models to produce physically-plausible or explainable results. For instance, Wi et al. [255] examine whether deep learning models can produce reliable future projections of streamflow under warming. It concludes that the hybrid LSTM using estimates of evapotranspiration from the process-based model as an additional input feature may support the use of LSTMs for hydrologic projections under climate change, but so may training LSTMs to a large, diverse set of watersheds. Using physical models (such as global climate models or global hydrological models) coupled with LSTM [256] is proposed for long-term streamflow projections. The LSTM model is considered a post-processor that aims to constrain (i.e., reduce the uncertainty of) streamflow simulations from the physics-based model. To enhance streamflow prediction in the weather research and forecasting hydrological modeling system (WRF-Hydro), a method called WRF-Hydro-LSTM [257] is proposed. It utilizes LSTM to predict the residual errors of WRF-Hydro.
@@
However, due to the presence of data errors in hydrological records, an unconstrained ML (such as LSTM) performs better than a PeML model (such as MC-LSTM) because of the ability of PDgML to learn and account for data errors [254]. In addition, three LSTM daily streamflow prediction models [258] are built and compared, including a LSTM post-processor trained on the United States National Water Model (NWM) outputs, a LSTM post-processor trained on the NWM outputs and atmospheric forcings, and a LSTM model trained only on atmospheric forcing. It concludes that a LSTM model trained only on atmospheric forcing outperforms the other two LSTMs in ungauged basins. Hybrid models may reduce the reliability of the model in some cases. Thus, PiML, PeML, and PaHL approaches for the reliability of rainfall-runoff forecasts require further development in the future.
@@
Missing Data and Ungauged Basins. See Fig. 5 (label 1.5). Hydrologic predictions at many watersheds, such as rural watersheds, as well as cold and ungauged regions, are important but also challenging due to data scarcity [278]. Thus, many PaML-based methods are developed for poorly gauged or ungauged basins. For example, geo-spatio-temporal mesoscale data and attention-based deep learning [259] are used to improve long-term streamflow prediction. This study demonstrates that the 3D-CNN–Transformer achieves the best performance compared with TD-CNN–Transformer, TD-CNN–LSTNet, and 3D-CNN–LSTNet, and the application of hydroclimate data with proposed integrated networks is particularly useful for poorly gauged basins. In addition, while deep neural networks demonstrated stellar performance for streamflow predictions, performance nonetheless declined for prediction in ungauged regions. In order to reduce such errors, an input-selection ensemble method [260] is proposed by leveraging the deep network’s flexibility to integrate satellite-based soil moisture products or daily flow distributions. Furthermore, hybrid models can be suitable alternatives to hydrological models, particularly in watersheds where there is a lack of measured data (e.g., climatic parameters, land cover-plant growth data, soil data, stream conditions, and properties of groundwater aquifers), provided that appropriate inputs (such as rainfall) are used. Mohammadi et al. [196] show that AI-based hybrid models have generally led to more accurate streamflow estimates compared with conceptual rainfall-runoff models in ungauged basins. To enhance out-of-distribution prediction in data-scarce basins, some serial methods for developing the hybrid LSTM [261, 262], i.e., using the physics-based hydrologic model simulation outputs as another set of inputs in the LSTM network, are proposed. Additionally, to reproduce long-duration daily streamflow in poorly gauged watersheds, the transfer learning technique is employed in a pre-trained SWAT-LSTM model [233].
@@
However, these PaML-based approaches are still only preliminary attempts. Deeply exploiting the transferability of ML models (such as domain adaptation methods [279], meta-learning methods [280], and pre-trained large models [foroumandichatgpt]) and the interpretability of physical models will effectively solve hydrological prediction problems in ungauged basins.

### v4

@@ 3 HydroPML: Process-based Hydrology in Physics-aware Machine Learning
As shown in Fig. 5, the proposed HydroPML includes rainfall-runoff hydrological process understanding and hydrodynamic process understanding.
@@ 3.1 Rainfall-runoff Process Understanding
Rainfall-runoff models are extensively utilized in hydrology to investigate hydrological processes and play a crucial role in water resources management, encompassing runoff prediction [203], flood prediction [204], and drought analysis [205]. These models have been developed for diverse applications, ranging from small catchments to global scales. The hydrological model can be classified into three broad groups: the empirical model, conceptual models, and physical-based models [2]. Detailed information regarding Physical Methods in Hydrological Process is available in Appendix A of Supplementary Material.
@@ 3.1.1 Rainfall-runoff Forecast Meets Physics-aware Machine Learning
(1.1) Short-term Forecasts LSTM [203, 206], physics-guided LSTM [32], GRUs [207], graph convolutional GRUs [208], graph convolutional RNN [209], stacked LSTM and bi-directional LSTM [210], ensemble model [211, 212], ConvLSTM [213], data-driven forecasting using only in-stream measurements [214] GW-PINN [215] Hybrid method combining hydrologic processes with LSTM [216], SWAT-LSTM [217], conceptual hydrological model by integrating NNs [218], urban rainfall-runoff modeling [219], Google’s operational flood forecasting system [220]
(1.2) Long-term Forecasts LSTM [10], Gaussian process regression [221], deep belief network [222], extreme learning machine [223], an encoder-decoder CNN algorithm [224], Bayesian ensemble learning by combining different ML methods [225] LSTM by considering conservation of mass into loss function [15] Entity-aware-LSTM [226], MC-LSTM [168], differentiable modeling [13], P-RNN layer [9] A hybrid monthly streamflow forecasting approach [194], ML is embedded into a lumped conceptual rainfall-runoff model [195], three hybridization approaches [197], differentiable parameter learning [227]
@@
(1.3) Spatial Variability Fully-distributed GNN [228], graph convolutional GRUs based model [208], graph convolutional RNN [209], ConvLSTM [213], graph-based reinforcement learning [229], data synergy [230], initializing a recurrent GNN by physical model [231] HydroNets [232], a differentiable, learnable physics-based routing model [233] Hybrid approach combining a spatially-distributed snow model with convLSTM [234]
@@
(1.4) Bias Reduction and Reliability Bias-corrected remote sensing products [235], directly measured datasets [214], ensemble methods [236, 237] MC-LSTM [168, 238], differentiable modeling [13] LSTM using an additional input feature from the process-based model [239], using physical models coupled with LSTM [240], WRF-Hydro-LSTM [241], LSTM daily streamflow prediction models [242]
@@
(1.5) Missing Data and Ungauged Basins Attention-based deep learning (such as 3D-CNN-Transformer) [243], input-selection ensemble method [244] AI-based hybrid models [196], serial hybrid LSTMs [245, 246], transfer learning in SWAT-LSTM [217].
@@
(1.6) Parameterization A machine learning approach [247], deep neural network [248], machine learning-based maps [249] A differentiable neural network [202], differentiable parameter learning [227]
@@
(1.7) Uncertainty Estimation in Rainfall-runoff Forecast Data-driven approach [250], Bayesian processor [251], hybrid ensemble and variational data assimilation framework [252], hierarchical model tree [253], probabilistic deep ensembles [254], LSTM-based approaches [255, 256], uncertainty estimation benchmarking procedure and baselines [257], analysis framework for sources of uncertainties [258] Using physical models coupled with LSTM [240], XAJ-MCQRNN [259], hybrid error correction model [260]
@@
Short-term Forecasts. See Fig. 5 (label 1.1). Short-term forecasts focus on outlook horizons ranging from hours to weeks. Innovative forecasting tools are transforming flood early warning mechanisms, and agricultural and hydropower management schemes. PaML offers robust tools to address the formidable challenges associated with the uncertainties, data dependencies, and intricacies inherent in short-term hydrological forecasting. First, the significant uncertainty and complexity inherent in hydrological processes and weather-climate factors affecting river basins have increasingly motivated researchers to adopt PDgML approaches. Specifically, the PDgML models based on LSTM [203, 206] are proposed to predict streamflow. They have been found to have greater accuracy in predicting hourly and daily streamflow than a classical neural network with back-propagation. A model [207] utilizing multiple GRUs is proposed for predicting streamflow up to 120 hours ahead. Granata et al. [211] demonstrate that their ensemble model, based on random forest and multilayer perceptron, outperforms bi-directional LSTM networks in predicting peak flow rates, while also achieving significantly shorter computation times. Wegayehu and Muluneh [210] compare various ML algorithms for one-step daily streamflow forecasting, showing that both MLP and GRU models perform better than stacked LSTM and bi-directional LSTM models to a similar extent. An ensemble model [212] incorporating nonlinear autoregressive networks, multilayer perceptrons, and random forests is proposed for short-term streamflow forecasting, using precipitation as the sole exogenous input, with a forecast horizon of up to 7 days. In addition, to model the spatial diversity of rivers, the graph convolutional GRUs-based model [208] is designed to predict streamflow for the next 36 hours at a sensor location by utilizing data from the upstream river network. Graph convolutional RNN [209] is developed for robust water demand forecasting. More detailed information about PDgML studies on short-term streamflow prediction can be found in [261]. Furthermore, to overcome data dependence, attempts have been made based on the PiML community, such as PINNs, for solving groundwater flow equations [215]. Because it is difficult to express the whole process of rainfall-runoff with specific equations, the current progress based on PiML technology is very slow. In addition, integrating hydrologic processes with LSTM [216] proves effective in assimilating recent streamflow observations, thereby enhancing near-term daily streamflow forecasts. A coupled soil and water assessment tool (SWAT)-LSTM approach [217] is developed to provide a potential shortcut for conducting daily streamflow simulations in both ungauged and poorly gauged watersheds. The conceptual hydrological model by integrating NNs [218] reveals that the hybrid approach outperforms both the original GR4J model and the single NN-based runoff prediction model in terms of prediction accuracy.
@@
In general, existing conceptual and physical models are widely used in operational forecasting within the time scale of hours to weeks. The development of PaML-based short-term forecast methods is not very mature [30]. However, PaML provides promising alternatives to traditional physics-based hydrological and inundation models for flood prediction due to its automation, efficiency, and flexibility. For example, the PDgML (ConvLSTM [213], data-driven forecasting using only in-stream measurements [214]) for extracting spatio-temporal features of hydrological information exceeds the recent models in predicting flood arrival time and peak discharge. PaHL-based methods (such as Google’s operational flood forecasting system [220] and urban rainfall-runoff modeling [219]) show potential for flood nowcasting.
@@
Long-term Forecasts. See Fig. 5 (label 1.2). Long-term forecasts focus on outlook horizons of sub-seasonal, year, to decade. The vast majority of ML predictions in the literature are PDgML methods, which are driven by observational or historical data. PDgML captures intricate non-linear patterns from data that would otherwise necessitate extrapolation, offering an alternative approach for long-term rainfall-runoff predictions, for example, LSTM [10], Gaussian process regression [221], deep belief network [222], extreme learning machine [223], an encoder-decoder CNN algorithm [224], and Bayesian ensemble learning by combining different ML methods [225]. These approaches demonstrate high accuracy in analyzing the relationship between rainfall and runoff, as well as the interaction between climate factors and rainfall. In order to alleviate the dependence on data, a physics-guided LSTM [15] that considers the conservation of mass into loss function is proposed. Digging deeper into physics (such as mass conservation and similarities) in the hydrological process, MC-LSTM [168] is introduced by modifying the LSTM to ensure mass input conservation. Entity-Aware-LSTM [226] learns catchment similarities as a feature layer within a DL model. In addition, based on the PeML technology, some NN modules with physical properties (such as differentiable modeling [13] and the P-RNN layer in the general physics-AI approach [9]) are proposed for rainfall-runoff forecasts. Furthermore, hybrid learning models (PaHL) can be suitable alternatives to hydrological models. Hybrid learning based on conceptual hydrological models tries to combine the strengths of hydrological models and data-driven methods. For example, a hybrid monthly streamflow forecasting approach [194] and three hybridization techniques [197] demonstrate that the hybrid model not only significantly surpasses conceptual hydrological models in precision but also outperforms purely ML models.
@@
However, in many hybrid modeling methods, variables generated by the conceptual model are utilized as inputs for a data-driven model, which generally leads to increased computation time. Additionally, these interactions are often overlooked because the two models are calibrated separately. Thus, embedded ML in a lumped conceptual rainfall-runoff model [195] is proposed by combining the two models and conducting their calibration jointly, and differentiable parameter learning [227] is presented to efficiently learn a hybrid model in an end-to-end manner.
@@
Spatial Variability. See Fig. 5 (label 1.3). Rainfall-runoff modeling is difficult due to the non-linear structure of the process, which involves much spatio-temporal variability. Spatial variability is critical for predictions in ungauged basins and for developing universal regional models. PaML (especially GNNs) provides the possibility of distributed modeling. Specifically, for PDgML, GNN [228] is used for fully-distributed rainfall-runoff modeling. The result shows that the GNN-based fully-distributed model [228] successfully represents the spatial information in predictions. A graph convolutional GRUs based model [208], ConvLSTM [213], and graph convolutional RNN [209] can also be developed to model spatial diversity of rivers. A recurrent GNN [231] is proposed to capture the interactions among multiple segments in the river network. In this study, a physics-informed model is employed to initialize the ML model and capture the underlying physics of streamflow. To mitigate the requirement for large training datasets, a real-time active learning approach is utilized within a graph-based reinforcement learning framework [229]. This method leverages spatial and temporal contextual information to select representative query samples. In addition, region-specific differences can be learned for the LSTM model by inputting sufficient heterogeneous data (data synergy) [230].
@@
Furthermore, spatial physical characteristics of rivers (such as river networks and hydrologic routing modules) can be encoded into NNs based on the PeML technology. For example, a hydrologic model called HydroNets [232] is built by leveraging river network structure. Specifically, each node in the graph represents a basin, and an edge direction corresponds to water flow from a sub-basin to its containing basin. A differentiable, learnable physics-based routing model [233] is proposed to constrain spatially-distributed parameterization. This marks the first instance of an interpretable, physics-based model being developed on the river network to infer spatially-distributed parameters. Finally, for the PaHL, the hybrid models combining conceptual models and data-driven models are often lumped, i.e., spatially averaged, at the catchment scale. To address this problem, a hybrid approach [234] is developed to predict streamflow. The hybrid approach combines a convLSTM with a spatially-distributed snow model to simulate the effect of surface and subsurface properties on streamflow. In the future, building a robust spatio-temporal representation for different watersheds based on the PaML community will be an ongoing direction for research.
@@
Bias Reduction and Reliability. See Fig. 5 (label 1.4). One important challenge of PaML-based rainfall-runoff forecasts is the ability of such models to produce reliable or physically-plausible results in different domains, especially extreme conditions such as floods and droughts. Specifically, reliable spatio-temporal remote sensing data and directly measured datasets can be used to reduce the bias of PDgML-based rainfall-runoff forecasting. For example, bias-corrected remote-sensing precipitation products are adopted as precipitation input of a spatio-temporal DL rainfall-runoff forecasting model [235], which can improve the model’s reliability for extreme flood forecasts. Directly measured datasets are utilized for streamflow estimation to develop a data-driven forecasting algorithm [214]. However, these bias-corrected data-based methods face challenges over small catchments due to uncertainties in remote sensing data and errors in hydrological models. Thus, some ensemble methods (such as resampling ensemble methods [236], which combine hydrometeorological modeling and LSTM [237]) are explored to improve the reliability and robustness of PDgML models. Furthermore, one emerging route for bias reduction and reliability is to employ PeML designs that explicitly observe the physical laws in the rainfall-runoff process, such as MC-LSTM [168] and differentiable modeling [13]. Another new development involves hybrid models designed to generate physically plausible or explainable results. For instance, Wi et al. [239] explore the capability of DL models to provide reliable future projections of streamflow under warming conditions. It concludes that integrating estimates of evapotranspiration from process-based models as additional input features can enhance the performance of LSTM models in hydrologic projections under climate change. Alternatively, training LSTMs on a diverse set of watersheds also proves effective. Another approach involves coupling physical models (such as global climate models or global hydrological models) with LSTM models [240] for long-term streamflow projections. Here, the LSTM acts as a post-processor aimed at constraining streamflow simulations derived from physics-based models to reduce uncertainty. To improve streamflow prediction within the Weather Research and Forecasting Hydrological Modeling System (WRF-Hydro), a method known as WRF-Hydro-LSTM [241] is proposed. This method utilizes LSTM models to predict residual errors within WRF-Hydro’s simulations.
@@
However, due to data errors in hydrological records, an unconstrained ML (such as LSTM) outperforms a PeML model (such as MC-LSTM) because of the ability of PDgML to learn and accommodate these data errors [238]. In addition, three LSTM daily streamflow prediction models [242] are built and compared, including a LSTM post-processor trained on outputs from the United States National Water Model (NWM), a LSTM post-processor trained on NWM outputs and atmospheric forcings, and a LSTM model trained solely on atmospheric forcings. It concludes that a LSTM model trained only on atmospheric forcing outperforms the other two LSTMs in ungauged basins. Hybrid models may reduce the reliability of the model in some cases. Thus, PiML, PeML, and PaHL approaches for the reliability of rainfall-runoff forecasts require further development in the future.
@@
Missing Data and Ungauged Basins. See Fig. 5 (label 1.5). Hydrologic predictions at many watersheds, such as rural watersheds, as well as cold and ungauged regions, are important but also challenging due to data scarcity [262]. Thus, many PaML-based methods are developed for poorly gauged or ungauged basins. For example, geo-spatio-temporal mesoscale data and attention-based deep learning [243] are used to improve long-term streamflow prediction. This study demonstrates the superior performance of the 3D-CNN–Transformer compared to TD-CNN–Transformer, TD-CNN–LSTNet, and 3D-CNN–LSTNet in hydroclimate data applications, particularly for poorly gauged basins. Despite the excellent performance of deep neural networks in streamflow predictions, their effectiveness declines in ungauged regions. To mitigate these errors, an input-selection ensemble method [244] is proposed, leveraging the flexibility of deep networks to integrate satellite-based soil moisture products or daily flow distributions. Furthermore, hybrid models present viable alternatives to hydrological models, especially in watersheds lacking measured data such as climatic parameters, land cover, soil data, and groundwater aquifer properties, provided that appropriate inputs like rainfall are available. Mohammadi et al. [196] demonstrate that AI-based hybrid models generally yield more accurate streamflow estimates than conceptual rainfall-runoff models in ungauged basins. To improve out-of-distribution prediction in data-scarce basins, several sequential methods have been proposed for developing hybrid LSTMs [245, 246]. These methods involve using outputs from physics-based hydrologic model simulations as additional inputs in LSTM networks. Additionally, to model long-duration daily streamflow in poorly gauged watersheds, transfer learning techniques are employed with a pre-trained SWAT-LSTM model [217].
@@
However, these PaML-based approaches are still only preliminary attempts. Deeply exploiting the transferability of ML models (such as domain adaptation methods [263], meta-learning methods [264]) and the interpretability of physical models will effectively solve hydrological prediction problems in ungauged basins. Furthermore, leveraging the increasing volume of simulation and observation data in hydrology, we can integrate large pre-trained models (e.g., generative models in ChatGPT [265], Transformer in ClimaX [266]) with hydrological knowledge to establish a foundational hydrology pre-training model. The pre-training large model can then be finely tuned to address diverse hydrological prediction tasks, including those in ungauged basins.